A car travellinga at 60km/h overtakes car 2 of speed 42km/h. Assuming each car to be5 m long,find the time taken during overtake and the total road distance used for the overtake.

Grade:12th pass

1 Answers

Aditya kashyap
6 years ago
Speed of car A is 60 km/hr or 16.67 m/s and of car B is 42 km/hr. Relative velocity of two car is C =60-42=18km/hr=5m/s. Each car is 5m long dist. For overtake is 10m. Time taken to overtake is t=10/5=2 s. And road dist. Covered is find by the speed of car A. As it overtake other car B . Dist=16.67×2=33.34m